# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v2

@@ 1. Introduction
However, prior works suggest that the queuing systems mix slower as one gets closer to the asymptotic regimes, for example, in the Heavy Traffic regime (Pang et al., 2007; Robert, 2003; Gamarnik and Goldberg, 2013). And so, we have a dilemma where the system has a larger steady-state approximation error if it is further away from the asymptotic regime, but mixes slower when it is closer to the asymptotic regime. And so, it is important to characterize the goodness of the steady-state analysis and asymptotic analysis to the finite-time queue length behavior. Therefore, the goal of this paper is to obtain queue length bounds in finite-time and pre-asymptotic regimes (i.e. with a finite number of servers) for the $M/M/n$ queue.
@@ 1.1. Contributions
where the $\square$ term is called the mixing rate of the system and $\chi$ is the square root of the Chi-square distance, which would be defined in Subsection 2.2. Our results are true for all time $t$ and for all $n$ (number of servers). Furthermore, our results are tight in the sense that as $n$ goes to infinity, our rate matches the mixing rate of $M/M/1$ and $M/M/\infty$, depending on the regime. We obtain finite-time convergence results for the following regimes:
@@
$\mathbb{E}\left[\#ofwaitingjobs\right]$ $O(1)$ $O(\sqrt{n})$ $O(n^{\alpha})$
Asymp. behavior Gaussian (Jhunjhunwala et al., 2023; Braverman et al., 2017) Exponential (Jhunjhunwala et al., 2023; Braverman et al., 2017)
@@
This work (van Leeuwaarden and Knessl, 2011; Gamarnik and Goldberg, 2013) This work This work, (Zeifman, 1991)
@@ 2.3. Main Results
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda\in\mathbb{Z}^{+}$ and the service rate $\mu$ and let the stationary distribution be $\nu_{\infty}$, we have:
@@ 2.3.1. Finite-time statistics
One can prove Corollary 4 by combining the Theorem 3 and Corollary 1 for $k=1$. We defer the details of the proof to Appendix C.1. Similar to the above corollary, we can obtain bounds on any moment $k\in\mathbb{Z}^{+}$ by Corollary 1 combined with a $2k$-moment bound on the steady-state distribution $\nu_{n}$. Moreover, one can go beyond the moments to also obtain tail bounds for the finite-time distribution which we present below.
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}.$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{1+\frac{1}{\delta}}\times e^{-\frac{x}{2(1+\delta)}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{4n^{1-2\alpha}}{7(1+\delta)}}+\left(10+\frac{2}{\delta}\right)n^{2\alpha-1}}\times e^{-\frac{x}{2(1+\delta)}}$
@@ 4.1. Regime 1: Super-Halfin-Whitt Regime
Observe that for $\alpha\geq 1$ or $n=1$, $K$ is a singleton as $2n^{1-\alpha}-1\leq 1$. In such a case, we simply apply Proposition 7 to obtain the desired mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ with no constant factor loss. Otherwise, if $\alpha\in(1/2,1)$ then $K$ is no longer a singleton. And so, we come to Step 2, where we wish to show that the system mixes quickly in the finite set $K$ by showing that the system admits a local Poincaré inequality inside the finite set. To this end, we apply the local canonical path method (described in Appendix B.1.3) to establish the local Poincaré constant $C_{L}=\Theta(n^{1-2\alpha})$ in Lemma 2, which means that $C_{L}$ goes to zero as $n$ goes to infinity and allows us to show rapid mixing in $K$.
@@
Note that we assume $n\geq 65$ only to obtain reasonable universal constants in the weighted-Poincaré constant. Obtaining an explicit bound for all $n\geq 1$ is possible, but doing so would result in unreasonably large universal constants. One can further improve such universal constants by increasing the threshold depending on the applications. The proof and the precise characterization of the constant $C_{b}$ is reserved in Appendix 5.1, which is done by a clever application of the local canonical path method. In the final step, we combine the results in Lemma 1 and Lemma 2 and apply Theorem 1 to obtain the results in Equation (6) and Equation (7) in Theorem 3 which gives us a convergence bound for $\alpha\in(1/2,\infty)$ that matches the limiting behavior at the asymptotics. The full detail of the proof is deferred to Appendix B.1.4.
@@ 4.2. Regime 2: Sub-Halfin-Whitt Regime
As we saw before in Proposition 3, the $M/M/n$ queue with $\alpha\in(0,1/2)$ behaves similarly to the $M/M/\infty$ queue. We use this connection to motivate our Lyapunov function. In the rest of the section, we first discuss the prove strategy for the $M/M/\infty$ queue, followed by that of the $M/M/n$ queue.
4.2.1. $M/M/\infty$
Consider the queue length $q$ of the $M/M/\infty$ queue with arrival rate $\lambda$ and service rate $\mu=1$. Unlike the $M/M/1$ queue, the dynamics of the diffusion analog of $(q-\lambda)/\sqrt{\lambda}$ is governed by the standard Ornstein-Uhlenbeck (OU) process. The Poincaré constant, and thus, the mixing rate of the OU process is equal to $1$ as a result of the Gaussian-Poincaré inequality (Bakry et al., 2013). One can show that $f(x)=x$ satisfy the drift equation (19) with exact equality for the OU-process with $\gamma=1$ and $b=0$, and so, it is a good choice for a Lyapunov function. One can also view $f(x)=x$ as the eigenfunction corresponding to the largest non-trivial eigenvalue $(=1)$ of the generator of the OU process, which is also the first non-constant Hermite polynomial. Hence, with the appropriate centering, we choose $V(q)=|q-\lambda|$ as our choice of the Lyapunov function to analyze the $M/M/\infty$ queue. In particular, analyzing the drift of $V(q)=|q-\lambda|$ allows us to prove Proposition 3. This result recovers the previous known mixing rate for the $M/M/\infty$ queue, see e.g.: (Chafaï, 2006). However, we present proof of this result using our Lypaunov-Poincaré methodology, which then motivates the proof for the $M/M/n$ queue in the sub-Halfin-Whitt regime.
4.2.2. $M/M/n$ in the Sub Halfin-Whitt regime
@@
The proof of this Lemma is deferred to Appendix B.3.1. Using the above lemma, the finite set is a singleton when $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, so we directly use Proposition 7 to infer the Poincaré constant, which in turn gives us the mixing rate. Observe that as $\gamma_{n}\rightarrow 1$ in the above lemma, we asymptotically match the mixing rate of the $M/M/\infty$ queue.
However, when $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, the finite set ceases to be a singleton. Nonetheless, $|K|=4$ is independent of $n$, and so we obtain a $\Omega(1/n)$ local Poincaré constant using the canonical path method since there are $\Theta(n)$ transitions per time unit. Although this is not enough to overcome the stitching error to get a mixing rate that approaches $1$, combining this local Poincaré constant with our drift lemma above using the Stitching Theorem (Theorem 1) is sufficient to obtain an $\Omega(1)$ mixing rate. We defer the full proof to Appendix B.3.
@@ A.1. Proof of Proposition 6
Let $\pi_{t}=\pi_{0}P_{t}$ and note that $\chi^{2}(\pi_{t},\nu)=\operatorname{Var}_{\nu}\left(\frac{\pi_{t}}{\nu}\right)=\operatorname{Var}_{\nu}\left(\frac{\pi_{0}P_{t}}{\nu}\right)$. Then, we have
$\frac{d}{dt}\chi^{2}(\pi_{t},\nu)=\frac{d}{dt}\operatorname{Var}_{\nu}\left(\frac{\pi_{0}P_{t}}{\nu}\right)=\frac{d}{dt}\left\lVert\frac{\pi_{0}P_{t}}{\nu}-1\right\rVert_{2,\nu}^{2}=2\left\langle\frac{d}{dt}\frac{\pi_{0}P_{t}}{\nu},\frac{\pi_{0}P_{t}}{\nu}-1\right\rangle_{\nu}.$
Since we have $\frac{d}{dt}\frac{\pi_{0}P_{t}}{\nu}=\lim_{\delta\rightarrow 0}\frac{\pi_{0}}{\nu}\left(\frac{P_{t+\delta}-P_{t}}{\delta}\right)=\frac{\pi_{0}P_{t}}{\nu}\lim_{\delta\rightarrow 0}\left(\frac{P_{\delta}-1}{\delta}\right)=\frac{\pi_{0}P_{t}\mathcal{L}}{\nu}$, this gives us
$\frac{d}{dt}\chi^{2}(\pi_{t},\nu)$ $=2\left\langle\frac{\pi_{0}P_{t}\mathcal{L}}{\nu},\frac{\pi_{0}P_{t}}{\nu}\right\rangle_{\nu}=2\left\langle\frac{\pi_{0}P_{t}}{\nu},\mathcal{L}\frac{\pi_{0}P_{t}}{\nu}\right\rangle_{\nu}\text{ from reversibility via the adjoint property,}$,
$=-2\mathcal{E}\left(\frac{\pi_{0}P_{t}}{\nu},\frac{\pi_{0}P_{t}}{\nu}\right)$
@@
where the inequality comes from the Poincaré inequality (16). Using Gronwall’s inequality, we have
@@ B.3.3. Proof of Proposition 5
We have that $V(q)\geq 1\,\forall q\leq n,q\neq\lambda_{n}$. For $q>n$, we choose $V(q)=\zeta e^{\theta\left(q-\lambda_{n}\right)}$ where $\zeta$ is a constant satisfies $|n-\lambda_{n}|=\zeta e^{\theta\left(n-\lambda_{n}\right)}$. Now, choose $\theta=\log\left(1+n^{\alpha_{n}-1}\right)$ such that $\delta>0$. Here, the term $\delta>0$ is added to ensure that the drift rate approaches to $1$ as in the proof of Lemma 3. Since $\theta>0$, we have that $V(q)\geq V(n)\geq 1\forall q\geq n$ and $V(q)\geq 1\forall q\leq n,q\neq\lambda_{n}$.
@@ B.5. Proof of Proposition 3
Here, we present the proof of $M/M/\infty$ convergence bound. As mentioned in the proof sketch (see Subsection 4.2.1), we want to choose $V(q)=|q-\lambda|$ as the Lyapunov function. Unfortunately, a direct application can only yield the right mixing bound when $\lambda\in\mathbb{Z}$, which will be shown as follows.
@@
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda\in\mathbb{Z}^{+}$ and the service rate $\mu$ such that $\lambda<\mu$ and let the stationary distribution be $\nu_{\infty}$, we have:
@@
Without the loss of generality, we assume that $\mu=1$. We will show that the Poincaré constant of the $M/M/\infty$ system is $1$. To this end, we will first analyze the drift of the system. Considering the Lyapunov function $V(q)=|q-\lambda|$, we want to show that $\mathcal{L}V\leq-V$ whenever $V$ takes some value outside of the singleton set $K=\{\lambda\}$ since $\lambda\in\mathbb{Z}^{+}$.
@@
$\mathcal{L}V(q)$ $=\lambda|q+1-\lambda|+q|q-1-\lambda|-(\lambda+q)|q-\lambda|$
$=-|q-\lambda|=-V(q).$
For $q\in K=\{\lambda\}$, we have:
$\mathcal{L}V(\lambda)$ $=\lambda|\lambda+1-\lambda|+\lambda|\lambda-1-\lambda|-(\lambda+\lambda)|\lambda-\lambda|$
$=2\lambda.$
Thus, we have shown that $M/M/\infty$ satisfies a Foster-Lyapunov condition with rate $-1$ when $\lambda$ is an integer. And so, we have that whenever $\lambda\in\mathbb{Z}$, we have a negative drift with rate $1$ outside the singleton set $K=\{\lambda\}$. From Proposition 6 and Proposition 7, the following holds for $\lambda\in\mathbb{Z}$
$\chi(\pi_{t,\infty},\nu_{\infty})\leq e^{-t}\chi(\pi_{0,\infty},\nu_{\infty}).$
Remark: The $e^{-\mu t}$ convergence rate matches the transient solution of $M/M/\infty$ for $\mu$ is the service rate of the system. Previously, there are multiple mixing proofs of $M/M/\infty$, including an entropic functional inequality proof (Chafaï, 2006). Our proof is the first Lyapunov-Poincaré proof for $M/M/\infty$. However, it is unfortunate that this proof only works for $\lambda\in\mathbb{Z}$, as otherwise we will not be able to obtain the mixing rate $1$ for $M/M/\infty$ due to the finite set in this case is no longer a singleton.

### v3

@@ 1. Introduction
However, prior works suggest that the queuing systems mix slower as one gets closer to the asymptotic regimes, for example, in the Heavy Traffic regime (Pang et al., 2007; Robert, 2003; Gamarnik and Goldberg, 2013). And so, we have a dilemma where the system has a larger steady-state approximation error if it is further away from the asymptotic regime, but mixes slower when it is closer to the asymptotic regime. And so, it is important to characterize the goodness of the steady-state analysis and asymptotic analysis to the finite-time queue length behavior. Therefore, the goal of this paper is to develop a theoretical framework to analyze the mixing time of reversible Markov chains, which entails Erlang-C systems. Having such a framework would allow us to obtain queue length bounds in finite-time and pre-asymptotic regimes (i.e. with a finite number of servers) for the $M/M/n$ queue and finite-time analysis for other various applications such as learning policies for queues (van Kempen et al., 2024).
@@ 1.1. Contributions
where the $\square$ term is called the mixing rate of the system and $\chi$ is the square root of the Chi-square distance, which would be defined in Subsection 2.2. Our results are true for all time $t$ and for all $n$ (number of servers). Furthermore, our results are tight in the sense that as $n$ goes to infinity, our mixing rate matches the mixing rate of $M/M/1$ and $M/M/\infty$, depending on the regime. In particular, for $\alpha\in(1/2,\infty)$, the system behavior resembles an $M/M/1$ system with service rate $n$ as most of the time all $n$ servers will be busy (i.e. the queue length is $n$ or above). On the other hand, for $\alpha\in(0,1/2)$, the system will have an idle server and resembles $M/M/\infty$. Therefore, the mixing time behavior of these systems will have
We obtain finite-time convergence results for the following regimes:
@@
$\mathbb{E}\left[[q-n]^{+}\right]$ $O(1)$ $O(\sqrt{n})$ $O(n^{\alpha})$
Tail behav. Gaussian Exponential
@@
Ours (Gamarnik and Goldberg, 2013) Ours Ours, (Zeifman, 1991)
@@ 2.3. Main Results
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda>0$ and the service rate $\mu>0$ such that $\lambda/\mu\in\mathbb{Z}^{+}$ and let the stationary distribution be $\nu_{\infty}$, we have:
@@ 2.3.1. Finite-time statistics
One can prove Corollary 4 by combining the Theorem 3 and Corollary 1 for $k=1$. We defer the details of the proof to Appendix C.1. Similar to the above corollary, we can obtain bounds on any moment $k\in\mathbb{Z}^{+}$ by Corollary 1 combined with a $2k$-moment bound on the steady-state distribution $\nu_{n}$. Moreover, in addition to the mean queue length bound, we can establish some tail-bound results to gain some understanding on the concetration of the finite-time behavior of the queue length distribution as follows.
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]\leq\left(1+\chi(\pi_{n,t},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}.$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon_{n}\left(q-n\right)>x\right]$ $\leq\left(1+e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{ex}e^{-\frac{x}{2}}$
@@
$\mathbb{P}_{\pi_{n,t}}\left[\varepsilon\left(q-n\right)>x\right]$ $\leq\left(1+e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{0.5e^{-\frac{2n^{1-2\alpha}}{7}}+6n^{2\alpha-1}}\times\sqrt{ex}e^{-\frac{x}{2}}$
@@ A CTMC admits a Poincaré constant $C_{P}$ if for all test function $f\in\ell_{2,\nu}$, we have
Comparing Corollary 2 with Proposition 7, one can see that when we try to ”stitch” negative drift outside of the finite set $K$ with the local mixing condition behavior inside the finite set, we incur a ”stitching error” $BC_{L}$ which hurts our Poincaré constant. Ideally, to get a tight mixing bound, we would like to make this stitching error term to be as small as possible or vanishes to 00 as $n$ goes to infinity. And so, it is crucial to choose the right Lyapunov function so that we get that right negative drift rate $\gamma$ while minimizing $B$, as well as developing methods to get tight local mixing constant $C_{L}$. We shall elaborate more on these steps in the following Section 4.
@@ 4.1. Regime 1: Super-Halfin-Whitt Regime
Observe that for $\alpha\geq 1$ or $n=1$, $K$ is a singleton as $2n^{1-\alpha}-1\leq 1$. In such a case, we simply apply Proposition 7 to obtain the desired mixing rate $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ without incurring any stitching error. For $\alpha\in(1/2,1)$, however, $K$ is no longer a singleton. And so, we come to Step 2, where we wish to show that the system mixes quickly in the finite set $K$ by showing that the system admits a local Poincaré inequality inside the finite set. To this end, we apply the local canonical path method (described in Lemma 1) to establish the local Poincaré constant $C_{L}=\Theta(n^{1-2\alpha})$ in Lemma 2, which means that $C_{L}$ goes to zero as $n$ goes to infinity and allows us to show rapid mixing in $K$.
@@
Note that we assume $n\geq 65$ only to obtain reasonable universal constants in the weighted-Poincaré constant. Obtaining an explicit bound for all $n\geq 1$ is possible, but doing so would result in unreasonably large universal constants. One can further improve such universal constants by increasing the threshold depending on the applications. The proof and the precise characterization of the constant $C_{b}$ is reserved in Subsection 5.1, which is done by a clever application of the local canonical path method. In the final step, we combine the results in Lemma 1 and Lemma 2 and apply Theorem 1 to obtain the results in Equation (6) and Equation (7) in Theorem 3 which gives us a convergence bound for $\alpha\in(1/2,\infty)$ that matches the limiting behavior at the asymptotics. The full detail of the proof is deferred to Appendix B.1.4.
@@ 4.2. Regime 2: Sub-Halfin-Whitt Regime
As we saw before in Proposition 3, the $M/M/n$ queue with $\alpha\in(0,1/2)$ behaves similarly to the $M/M/\infty$ queue. We use this connection to motivate our Lyapunov function. In the rest of the section, we first discuss the proof strategy for the $M/M/\infty$ queue, followed by that of the $M/M/n$ queue.
4.2.1. $M/M/\infty$ system
For an $M/M/\infty$ system with arrival rate $\lambda$ and service rate $\mu$, it is known that the system is a discrete space diffusion analogue of the Ornstein–Uhlenbeck (OU) process (Chafaï, 2006). For the latter process, it is also known that it admits the Gaussian Poincaré inequality with constant $1$, which implies unit mixing rate for the OU process and also suggests that the mixing rate for $M/M/\infty$ is $\mu$. To construct an appropriate Lyapunov function, one can look at the eigenfunction that corresponds to the correct mixing rate, which is also the smallest positive real number such that the spectral measure has a jump. For $M/M/\infty$, this eigenfunction is the first non-trivial Charlier polynomial $x-\frac{\lambda}{\mu}$ (de la Iglesia, 2021). Hence, with an appropriate centering, we choose $V(q)=\left|q-\frac{\lambda}{\mu}\right|$ as the Lyapunov function so that it captures the correct eigenvalue equation $\mathcal{L}V(q)=-\mu V(q)$ for all $q$ outside of some finite set $K$ and satisfies the non-negativity condition for Lyapunov functions in Assumption 2.
4.2.2. $M/M/n$ system in the Sub Halfin-Whitt regime
@@
The proof of this Lemma is deferred to Appendix B.3.1. Using the above lemma, the finite set is a singleton when $\lambda_{n}\in\mathbb{Z}_{\geq 0}$, so we directly use Proposition 7 to infer the Poincaré constant, which in turn gives us the mixing rate. Observe that as $\gamma_{n}\rightarrow 1$ in the above lemma, we asymptotically match the mixing rate of the $M/M/\infty$ queue. However, when $\lambda_{n}\notin\mathbb{Z}_{\geq 0}$, the finite set ceases to be a singleton. Nonetheless, $|K|=4$ is independent of $n$, and so we obtain a $\Omega(1/n)$ local Poincaré constant using the canonical path method since there are $\Theta(n)$ transitions per time unit. Although this is not enough to overcome the stitching error to get a mixing rate that approaches $1$, combining this local Poincaré constant with our drift lemma above using the Stitching Theorem (Theorem 1) is sufficient to obtain an $\Omega(1)$ mixing rate. We defer the full proof to Appendix B.3.
@@ A.1. Proof of Proposition 6
Note that $\chi^{2}(\pi_{t},\nu)=\operatorname{Var}_{\nu}\left(\frac{\pi_{t}}{\nu}\right)=\operatorname{Var}_{\nu}\left(P_{t}\frac{\pi_{0}}{\nu}\right)$, we have
$\frac{d}{dt}\chi^{2}(\pi_{t},\nu)=\frac{d}{dt}\operatorname{Var}_{\nu}\left(P_{t}\frac{\pi_{0}}{\nu}\right)=\frac{d}{dt}\left\lVert P_{t}\frac{\pi_{0}}{\nu}-1\right\rVert_{2,\nu}^{2}=2\left\langle\frac{d}{dt}P_{t}\frac{\pi_{0}}{\nu},P_{t}\frac{\pi_{0}}{\nu}-1\right\rangle_{\nu}.$
Since we have $\frac{d}{dt}P_{t}\frac{\pi_{0}}{\nu}=\lim_{\delta\rightarrow 0}\left(\frac{P_{t+\delta}-P_{t}}{\delta}\right)\frac{\pi_{0}}{\nu}=\lim_{\delta\rightarrow 0}\left(\frac{P_{\delta}-1}{\delta}\right)P_{t}\frac{\pi_{0}}{\nu}=\mathcal{L}P_{t}\frac{\pi_{0}}{\nu}$ (due to the fact that $P_{t+s}=P_{t}P_{s}$ for any $t,s$), this gives us
$\frac{d}{dt}\chi^{2}(\pi_{t},\nu)$ $=2\left\langle\mathcal{L}P_{t}\frac{\pi_{0}}{\nu},P_{t}\frac{\pi_{0}}{\nu}\right\rangle_{\nu}=2\left\langle P_{t}\frac{\pi_{0}}{\nu},\mathcal{L}P_{t}\frac{\pi_{0}}{\nu}\right\rangle_{\nu}\text{ from reversibility via the adjoint property,}$,
$=-2\mathcal{E}\left(P_{t}\frac{\pi_{0}}{\nu},P_{t}\frac{\pi_{0}}{\nu}\right)$
@@
where the inequality comes from the Poincaré inequality (16). From Gronwall’s inequality, we have
@@ B.3.3. Proof of Proposition 5
We have that $V(q)\geq 1\,\forall q\leq n,q\neq\lambda_{n}$. For $q>n$, we choose $V(q)=\zeta e^{\theta\left(q-\lambda_{n}\right)}$ where $\zeta$ is a constant satisfies $|n-\lambda_{n}|=\zeta e^{\theta\left(n-\lambda_{n}\right)}$. Now, choose $\theta=\log\left(1+n^{\alpha_{n}-1}\right)$, we have $e^{\theta}-1=n^{\alpha_{n}-1}$. Since $\theta>0$, we have that $V(q)\geq V(n)\geq 1\forall q\geq n$ and $V(q)\geq 1\forall q\leq n,q\neq\lambda_{n}$.
@@ B.5. Proof of Proposition 3
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda>0$ and the service rate $\mu>0$ such that $\lambda/\mu\in\mathbb{Z}^{+}$ and let the stationary distribution be $\nu_{\infty}$, we have:
@@
Considering the Lyapunov function $V(q)=|q-\lambda/\mu|$, we want to show that $\mathcal{L}V\leq-\mu V$ whenever $V$ takes some value outside of the singleton set $K=\{\lambda/\mu\}$ since $\lambda/\mu\in\mathbb{Z}^{+}$.
@@
$\mathcal{L}V(q)$ $=\lambda|q+1-\lambda/\mu|+q\mu|q-1-\lambda/\mu|-(\lambda+q\mu)|q-\lambda|=-\mu|q-\lambda/\mu|=-\mu V(q).$
For $q\in K=\{\lambda/\mu\}$, we have:
$\mathcal{L}V(\lambda)$ $=\lambda|\lambda/\mu+1-\lambda/\mu|+\lambda/\mu|\lambda/\mu-1-\lambda/\mu|-(\lambda+\lambda/\mu)|\lambda/\mu-\lambda/\mu|=\lambda+\lambda/\mu.$
Thus, we have shown that $M/M/\infty$ satisfies a Foster-Lyapunov condition with rate $-\mu$ outside the singleton set $K=\{\lambda/\mu\}$ when $\lambda/\mu$ is an integer. From Proposition 6 and Proposition 7, the following holds for $\lambda/\mu\in\mathbb{Z}$
$\chi(\pi_{t,\infty},\nu_{\infty})\leq e^{-\mu t}\chi(\pi_{0,\infty},\nu_{\infty}).$
Hence proved. ∎
Remark: The $e^{-\mu t}$ convergence rate matches the transient solution of $M/M/\infty$ for $\mu$ is the service rate of the system. Previously, there are multiple mixing proofs of $M/M/\infty$, including an entropic functional inequality proof (Chafaï, 2006). Our proof is the first Lyapunov-Poincaré proof for $M/M/\infty$. However, it is unfortunate that this proof only works for $\lambda/\mu\in\mathbb{Z}$, as otherwise we will not be able to obtain the mixing rate $\mu$ for $M/M/\infty$ due to the finite set in this case is no longer a singleton.